# A Prototype Hybrid Mode Cavity for Heterodyne Axion Detection

### v1

@@ 3.1 Hybrid Mode Concept
If we directly took the two $\text{HE}_{11}$ modes to be the loaded and signal mode, then $\tilde{\mathbf{E}}_{1}^{*}$ would be parallel to $\tilde{\mathbf{B}}_{0}$ but offset by $\lambda/4$, leading to zero signal form factor. To resolve this problem, fins of depth $\lambda/4$ are added to the endplates. As shown in Fig. 2a, the two fins are oriented orthogonally, so that each set reflects one polarization while transmitting the other. Since the fields of the modes are almost completely transverse, this shift leads to a near-perfect alignment of $\tilde{\mathbf{E}}_{1}^{*}$ and $\tilde{\mathbf{B}}_{0}$ within the cavity, except in the endplate regions, as shown in Fig. 2c. Since the side walls carry $21$ corrugations spaced by $\lambda/4$, we roughly estimate $C_{\text{sig}}\approx 19/21\approx 0.9$.
@@ 3.2 Optimized Design
The profiles of the hybrid modes are shown in Figs. 3b and 3c. The “fixed mode” has a frequency of $2.8565\,\mathrm{GHz}$ and an intrinsic quality factor of $Q=7.0\times 10^{4}$ for a cavity entirely made of room temperature copper. The “tunable mode” has a frequency set by the location of the tuning membrane, as will be discussed below. It has a slightly lower expected quality factor of $Q=6.3\times 10^{4}$ due to the presence of the aluminum tuning plate. We did not calculate an exact value for the signal form factor, as it involves a volume integral which is nontrivial to perform within ACE3P; however, since the modes are qualitatively similar to the cylindrical case, we estimate $C_{\text{sig}}\simeq 0.9$.
@@ 3.3 Coupling and Frequency Tuning
Table 1: Maximum magnitudes of the field profiles within the bulk of the cavity and on its surface, in units of $\text{m}^{-3/2}$. The profiles are normalized so that $\int_{V}|\tilde{E}_{i}|^{2}=\int_{V}|\tilde{B}_{i}|^{2}=1$, so that their rms value is $1/\sqrt{V}=2.9\,\mathrm{m}^{-3/2}$. The fields are largest on the endplate fins, and further enhanced when a large coupling hole is present, which is required for a non-superconducting setup.
@@
The couplers are designed for critical coupling in the prototype cavity, and therefore require an external quality factor of approximately $7.0\times 10^{4}$. To minimize mode perturbations and field enhancements of the overmoded $\text{HE}_{11}$ mode due to the coupling hole, a quarter-wave transformer is incorporated into the coupler waveguide. This transformer enhances the coupling efficiency, allowing for a smaller coupling hole with less impact on the mode structure. The coupler design used in the prototype cavity is shown at left in Fig. 4.
@@ 3.4 Cross-Coupling Suppression
Figure 5: Simulations of cross-coupling suppression. Imperfections of the cavity geometry lead to a larger $\epsilon$, which can be suppressed by rotating an endplate. Dots show simulated results, and the dashed curve is an interpolation. To demonstrate the point, we consider simulated imperfections of the cavity which are much more severe than what would actually occur. (a) Half of the cavity is twisted by $0.1^{\circ}$. Rotating the endplate by a similar amount restores a low value of the cross-coupling, $\epsilon\lesssim 10^{-4}$. (b) The cavity is skewed, so that one diagonal is $0.8\,\mathrm{mm}$ longer than the other. Rotating the endplate restores $\epsilon\lesssim 10^{-3}$.
In the ideal cavity geometry the cross-coupling vanishes exactly, and this remains true when the cavity is tuned in an azimuthally uniform way. However, in practice cross-coupling can be introduced by fabrication imperfections and misalignments. Some of these effects can be mitigated by incorporating tuning mechanisms into the cavity, e.g. by rotating the tunable endplate about the cavity axis.
To check the feasibility of this concept, we used ACE3P to compute the cross-coupling at the fixed mode frequency, for several choices of the cavity geometry. First, we considered a situation where half of the cavity is twisted along the cavity axis by $0.1^{\circ}$ relative to the other half. This is a much larger imperfection than one would expect in the fabrication process, and it degrades the mode isolation to $\epsilon\sim 10^{-1}$. However, as shown in Fig. 5a, rotating the endplate can restore the isolation to $\epsilon\lesssim 10^{-4}$.
This result might not be particularly surprising, as both the simulated imperfection and the endplate adjustment involve rotation about the same axis. As a more extreme example, we considered a skew of the entire cavity, so that its diagonals are lengthened and shortened by $0.4\,\mathrm{mm}$. Again, this is a substantially larger change than one would expect from fabrication, and it degrades the mode isolation to $\epsilon\sim 10^{-1}$. As shown in Fig. 5b, by rotating the endplate, the simulated mode isolation can be recovered to $\epsilon\sim 10^{-3}$.
@@ 4.1 Cavity Assembly
The fixed endplate is directly mounted on the cavity barrel, but the tunable endplate is mounted against two jack screws with a range of motion of ±$\pm 250\,\mu\mathrm{m}$, whose heights can be used to perform a roll angle adjustment, i.e. to rotate this endplate about the cavity axis. Once the desired orientation is found, the endplate is tightened against the central barrel using wide flange screws. One could also adjust the yaw angle of the tunable endplate by controlling the differential height of additional jack screws, but this was not pursued in the current prototype.
@@ 4.2 Tuning Mechanism
As shown in Fig. 1b, the tunable endplate contains a large square opening, into which is placed an aluminum tuning plate with a thin circular membrane of $1\,\mathrm{mm}$ thickness, as shown in Fig. 7a. The membrane was designed so that it could be deformed through a total range of $2\,\mathrm{mm}$, to yield the desired $\sim 10\,\mathrm{MHz}$ tuning range discussed in Sec. 3.3. The membrane is initially bowed outward, and in its innermost position, it is roughly flush with the wall of the tunable endplate, so that the tunable mode’s frequency is approximately equal to that of the fixed mode.
@@ 5 Characterization of the Cavity
Figure 10: Measurements of the microwave scattering parameters. (a) Demonstration of the tuning range. As the tuning screws were rotated, the frequency of the tunable mode increased by about $4\,\mathrm{MHz}$, while the fixed mode frequency remained the same. (b) Demonstration of decoupling. Top: at a point in the middle of the tuning range, the cross-coupling $\epsilon^{2}=|S_{\mathrm{tf}}|^{2}$ was originally $\sim 10^{-7}$. By rotating the tunable endplate, it was reduced to $\sim 10^{-8}$, the noise floor of the analyzer. Bottom: the decoupling procedure was repeated at another frequency, with similar results.
@@ 5.1 Mode Properties and Frequency Tuning
[…]
@@ 5.2 Suppression of Cross-Coupling
This measurement demonstrates that a cross-coupling $\epsilon\sim 10^{-4}$ can be achieved in our prototype cavity, corresponding to 80 dB of noise reduction. Even further reduction may be possible, but would not be visible in our measurement setup given the noise floor of the network analyzer.
@@ 6.1 Related Work and Future Directions
The modern wave of interest in heterodyne axion detection began in 2020 [16, 17, 18], and for the past several years, several experimental groups have been developing prototypes. The SHADE collaboration at Fermilab’s SQMS has characterized noise in a driven 9-cell superconducting cavity [37], and developed single cell cavities with a small, tunable splitting between the $\mathrm{TM}_{020}$ and $\mathrm{TE}_{011}$ mode [38]. The group also performed a sensitive search for dark photon dark matter using an unloaded superconducting cavity with fixed frequency [39]. The SHANHE collaboration at Peking University extended the latter result by demonstrating mechanical frequency tuning at cryogenic temperatures [40, 41], and is developing superconducting cavities optimized for heterodyne axion detection. In addition, an Australian group operated a heterodyne search with a small cylindrical copper cavity across a $\sim 10\%$ tuning range in $m_{a}$ [42]. Finally, the CERN Quantum Technology Initiative, in affiliation with the Physics Beyond Colliders group, has recently begun developing superconducting cavities for heterodyne axion detection with non-mechanical tuning mechanisms [43].
@@ 6.2 Outlook for a Superconducting Prototype
Our prototype has been designed from the ground up to be optimized for heterodyne axion detection. Its high signal form factor $C_{\text{sig}}\simeq 0.9$ compares favorably with the values $C_{\text{sig}}\simeq 0.5$ and $C_{\text{sig}}\simeq 0.2$ achievable in more traditional cylindrical or accelerator cavity designs, respectively, and could yield a substantially higher signal power $P_{\text{sig}}\propto C_{\text{sig}}^{2}$. Even more radical designs with toroidal cavity geometries have been proposed [17, 44], though it is less clear how to fabricate and tune them.
@@
Leakage noise measurements could be carried out in an existing cryogenic test facility, and would inform other aspects of the superconducting design. For instance, typical niobium cavities have a thickness of a few mm, while our current design has endplates of $20\,\mathrm{mm}$ thickness. Thicker walls would greatly reduce vibrations, but may make thermal management more difficult. Moreover, stiffening the tuning membrane might lower the tuning range. The appropriate tradeoff can be determined by performing in situ measurements.
@@
To estimate the potential sensitivity, we note that for a thermal noise limited heterodyne axion experiment with amplifier noise at the standard quantum limit, the signal to noise ratio is [16]
@@
$\text{SNR}\sim\frac{\rho_{\mathrm{DM}}\,V}{m_{a}\,\omega_{1}}\,(g_{a\gamma\gamma}\,C_{\text{sig}}\,B_{0})^{2}\sqrt{\frac{Q_{a}\,Q_{\text{int}}\,t_{e}}{T}}.$ (12)
@@
This holds up to an order-one factor which depends on details such as the precise shape of the axion frequency spectrum, and assumes the frequency scan is performed uniformly in $\log(m_{a})$. It also assumes that the readout waveguide is overcoupled by a factor of $T/\omega_{1}$, which optimizes the scan rate. For the dark matter, we take $\rho_{\mathrm{DM}}=0.4\,\text{G}\text{e\kern-0.6458ptV}/\mathrm{cm}^{3}$ and $Q_{a}=10^{6}$.
@@
For the cavity, we take the prototype values $C_{\text{sig}}=0.9$, $V=0.12\,\mathrm{m}^{3}$ and $\omega_{1}=2\pi\,(2.85\,\mathrm{GHz})$, and assume cooling to $T=2\,\mathrm{K}$. For these parameters, the overcoupling factor is $T/\omega_{1}\sim 15$. In our prototype design, our simulations indicated that a tuning range of $10\,\mathrm{MHz}$ is both mechanically feasible and preserves the cavity mode properties. Only about half this tuning range was realized in the prototype cavity, since we could only push the tuning membrane inward. We assume that the motorized tuning mechanism for the superconducting cavity can both push and pull, realizing the full potential tuning range. It is simplest to estimate the sensitivity at high frequencies, where leakage noise is automatically suppressed, so we suppose the mode frequency difference is scanned from $1\,\mathrm{MHz}$ to $11\,\mathrm{MHz}$.
@@
As for the intrinsic quality factor $Q_{\text{int}}$, we have seen that a pure copper prototype cavity would have $Q_{\text{int}}=7\times 10^{4}$ and a surface resistance $R_{s}=\sqrt{\omega_{0}/2\sigma}=1.4\times 10^{-2}\,\Omega$, where $\sigma$ is the standard conductivity of room temperature copper. The surface resistance of niobium depends on frequency, temperature, and the manufacturing process. It can be decomposed as $R_{s}=R_{\text{BCS}}(f,T)+R_{\text{res}}(f)$, where the BCS contribution is proportional to $f^{2}$ and the residual resistance is due to impurities, defects, and trapped magnetic flux [45]. For the LCLS-II project, commercial vendors use nitrogen doping to reliably produce superconducting cavities with $R_{\text{BCS}}(1.3\,\mathrm{GHz},2\,\mathrm{K})=4.5\,\mathrm{n}\Omega$ and $R_{\text{res}}=2.5\,\mathrm{n}\Omega$ [46]. We assume the residual resistance is dominated by trapped flux, so that $R_{\text{res}}\propto f^{1/2}$ [45, 47]. This yields $R_{s}=25\,\mathrm{n}\Omega$ at the prototype cavity frequency, which by $Q_{\text{int}}\propto 1/R_{s}$ gives a projected quality factor $Q_{\text{int}}=3.9\times 10^{10}$.
@@
In Fig. 11, we show the projected reach of a superconducting version of our prototype at $\text{SNR}=2$. The “test run” assumes a total scan time of 1 week and an input power $P_{\mathrm{in}}=0.1\,\mathrm{W}$, which is within the reach of a standard low-noise oscillator. In this regime leakage noise would almost certainly be negligible, given the low values of $\epsilon$ and $\eta$ in our design, the small input power, and the $\gtrsim\mathrm{MHz}$ axion frequency. Note that there is sufficient time for the signal to ring up at each step, since the signal mode is overcoupled. The “full run” curve assumes the same total scan time, but a loaded mode with typical magnetic field $B_{0}=0.03\,\mathrm{T}$. Referring to Table 1, this corresponds to a maximum magnetic field of $0.2\,\mathrm{T}$ at the cavity walls, due to the higher fields at the endplate fins. Maintaining this field requires an input power $P_{\mathrm{in}}=\omega_{0}B_{0}^{2}V/2Q_{\mathrm{int}}=20\,\mathrm{W}$, which is well within the capacity of a vertical test stand. Finally, the “double scale” curve assumes a scan time of 1 year and a direct scale-up of the cavity with no other changes. This automatically halves the mode frequencies, multiplies the volume by $8$, and increases the quality factor to $2.4\times 10^{11}$.

### v2

@@ 3.1 Hybrid Mode Concept
If we directly took the two $\text{HE}_{11}$ modes to be the loaded and signal mode, then $\tilde{\mathbf{E}}_{1}^{*}$ would be parallel to $\tilde{\mathbf{B}}_{0}$ but offset by $\lambda/4$, leading to a cancellation in the signal form factor in Eq. (5). To resolve this problem, fins of depth $\lambda/4$ can be added to the endplates. As shown in Fig. 2a, each set of fins reflects the mode whose electric field is polarized parallel to the fins, while allowing the other mode to propagate through and reflect off the end of the cavity. By orienting the fins orthogonally, one mode is confined between the left end of the cavity and the left edge of the right set of fins, while the other is confined between the right edge of the left set of fins and the right end of the cavity, as shown in Fig. 2c.
As a result, though both modes continue to see a cavity of equal effective length, the fins offset their profiles by a compensating $\lambda/4$, so that now $\tilde{\mathbf{E}}_{1}^{*}$ and $\tilde{\mathbf{B}}_{0}$ have the same phase within the cavity. Furthermore, since these are overmoded balanced hybrid modes, $\tilde{\mathbf{E}}_{1}^{*}$ and $\tilde{\mathbf{B}}_{0}$ are approximately parallel to each other, with the same spatial distribution, maximizing the signal form factor. Note that within each set of fins, one mode is present (and roughly doubled in magnitude) while the other is exponentially suppressed. Conservatively assuming that these regions contribute zero to the signal form factor, we roughly estimate $C_{\text{sig}}\approx 19/21\approx 0.9$.
@@ 3.2 Optimized Design
geometry factor $G^{(s)}\,(\Omega)$ participation ratio $p_{s}$
surface tunable fixed tunable fixed
left side plate $1.40\times 10^{5}$ $7.98\times 10^{4}$ $0.007$ $0.013$
right side plate $1.48\times 10^{5}$ $8.04\times 10^{4}$ $0.007$ $0.013$
bottom side plate $7.57\times 10^{4}$ $9.41\times 10^{4}$ $0.013$ $0.011$
top side plate $7.54\times 10^{4}$ $9.35\times 10^{4}$ $0.013$ $0.011$
fixed endplate $5.28\times 10^{3}$ $1.32\times 10^{3}$ $0.186$ $0.764$
tunable endplate $1.27\times 10^{3}$ $5.33\times 10^{3}$ $0.774$ $0.190$
full cavity $9.83\times 10^{2}$ $01\times 10^{3} 11$ 11
Table 1: Contributions of each side of the cavity to the mode loss. Under our normalization $\int_{V}|\tilde{B}_{i}|^{2}=1$, the geometry factor for a given surface satisfies $1/G^{(s)}_{i}=\int_{s}|\tilde{B}_{i}|^{2}\,dS/(\mu_{0}\omega)$, where $\omega$ is the mode angular frequency. The participation ratio $p_{s}\propto 1/G^{(s)}_{i}$ gives its fractional contribution to the total loss.
@@
The profiles of the hybrid modes are shown in Figs. 3b and 3c. Table 1 shows how each side of the cavity contributes to the loss; as expected for hybrid modes, losses at the side plates are very small, so that losses at the endplates dominate. For a cavity entirely made of room temperature copper, the “fixed mode” has an intrinsic quality factor of $Q=7.0\times 10^{4}$, while the “tunable mode” has a slightly lower expected quality factor of $Q=6.3\times 10^{4}$ due to the presence of the aluminum tuning plate, to be discussed below. The fixed mode has a frequency of $2.8565\,\mathrm{GHz}$, while the frequency of the tunable mode depends on the location of the tuning membrane.
We did not calculate an exact value for the signal form factor, as it involves a volume integral which is nontrivial to perform within ACE3P; however, since the modes are qualitatively similar to the cylindrical case, we estimate $C_{\text{sig}}\simeq 0.9$.
@@ 3.3 Coupling and Frequency Tuning
Table 2: Maximum magnitudes of the field profiles within the bulk of the cavity and on its surface, in units of $\text{m}^{-3/2}$. The profiles are normalized so that $\int_{V}|\tilde{E}_{i}|^{2}=\int_{V}|\tilde{B}_{i}|^{2}=1$, so that their rms value is $1/\sqrt{V}=2.9\,\mathrm{m}^{-3/2}$. The fields are largest on the endplate fins, as shown in red in Fig. 3. They are also further enhanced when a large coupling hole is present, which is only required for a non-superconducting setup.
@@
The couplers are designed for critical coupling in the prototype cavity, i.e. RF coupling $\beta_{i}=1$, and therefore require an external quality factor of approximately $7.0\times 10^{4}$. To minimize mode perturbations and field enhancements of the overmoded $\text{HE}_{11}$ mode due to the coupling hole, a quarter-wave transformer is incorporated into the coupler waveguide. This transformer enhances the coupling efficiency, allowing for a smaller coupling hole with less impact on the mode structure. The coupler design used in the prototype cavity is shown at left in Fig. 4.
@@ 3.4 Cross-Coupling Suppression
Figure 5: Simulations of cross-coupling suppression. Imperfections of the cavity geometry lead to a larger $\chi_{\mathrm{d}}$ and $\chi_{\mathrm{r}}$, which can be simultaneously suppressed by rotating an endplate. Dots show simulated results, and the dashed curve is an interpolation. To demonstrate the point, we consider simulated imperfections of the cavity which are much more severe than what would actually occur. (a) Half of the cavity is twisted by $0.1^{\circ}$. Rotating the endplate by a similar amount restores a low value of the cross-coupling, $\chi\lesssim 10^{-4}$. (b) The cavity is skewed, so that one diagonal is $0.8\,\mathrm{mm}$ longer than the other. Rotating the endplate restores $\chi\lesssim 10^{-3}$.
The cavity with two couplers is a two-port network, with each port dominantly coupled to one of the orthogonal $\mathrm{HE}_{11}$ modes. As described in Eq. (8), the scattering parameter $S_{\mathrm{rd}}$ is used to quantify the cross-coupling. In the ideal cavity geometry the cross-coupling vanishes exactly, and this remains true when the cavity is tuned in an azimuthally uniform way. However, in practice cross-coupling can be introduced by fabrication imperfections and misalignments. Some of these effects can be mitigated by incorporating tuning mechanisms into the cavity, e.g. by rotating the tunable endplate about the cavity axis.
To check the feasibility of this concept, we used ACE3P to compute the cross-coupling at the fixed mode frequency, for several choices of the cavity geometry. First, we considered a situation where half of the cavity is twisted along the cavity axis by $0.1^{\circ}$ relative to the other half. This is a much larger imperfection than one would expect in the fabrication process, and it degrades the mode isolation to $\chi\sim 10^{-1}$. However, as shown in Fig. 5a, rotating the endplate can restore the isolation to $\chi\lesssim 10^{-4}$.
This result might not be particularly surprising, as both the simulated imperfection and the endplate adjustment involve rotation about the same axis. As a more extreme example, we considered a skew of the entire cavity, so that its diagonals are lengthened and shortened by $0.4\,\mathrm{mm}$. Again, this is a substantially larger change than one would expect from fabrication, and it degrades the mode isolation to $\chi\sim 10^{-1}$. As shown in Fig. 5b, by rotating the endplate, the simulated mode isolation can be recovered to $\chi\sim 10^{-3}$.
@@ 4.1 Cavity Assembly
The fixed endplate is directly mounted on the cavity barrel, but the tunable endplate is mounted against two jack screws with a range of motion of $\pm 250\,\mu\mathrm{m}$, whose heights can be used to perform a roll angle adjustment, i.e. to rotate this endplate about the cavity axis. Once the desired orientation is found, the endplate is tightened against the central barrel using wide flange screws. One could also adjust the yaw angle of the tunable endplate by controlling the differential height of additional jack screws, but this was not pursued in the current prototype.
@@ 4.2 Tuning Mechanism
As shown in Fig. 1b, the tunable endplate contains a large square opening, into which is placed an aluminum tuning plate with a thin circular membrane of $1\,\mathrm{mm}$ thickness, as shown in Fig. 7a. The membrane was designed so that it could be deformed through a total range of $2\,\mathrm{mm}$, to yield the desired $\sim 10\,\mathrm{MHz}$ tuning range discussed in Sec. 3.3. The membrane is initially bowed outward, and in its innermost position, it is roughly flush with the wall of the tunable endplate, so that the tunable mode frequency is approximately equal to that of the fixed mode.
@@ 5 Characterization of the Cavity
Figure 10: Measurements of the microwave scattering parameters. (a) Demonstration of the tuning range. As the tuning screws were rotated, the frequency of the tunable $\mathrm{HE}_{11}$ mode increased by about $4\,\mathrm{MHz}$, while the frequency of the fixed $\mathrm{HE}_{11}$ remained the same. (b) Demonstration of decoupling. Top: at a point in the middle of the tuning range, the cross-coupling $|S_{\mathrm{tf}}|^{2}$ was originally $\sim 10^{-7}$. By rotating the tunable endplate, it was reduced to $\sim 10^{-8}$, the noise floor of the analyzer. Bottom: the decoupling procedure was repeated at another frequency, with similar results.
@@ 5.1 Mode Properties and Frequency Tuning
By measuring $S_{\mathrm{ff}}(f)$, shown in blue in Fig. 10a, we inferred that the fixed mode had frequency $2.8556\,\text{GHz}$. Using the $Q$-circle method, we found that its intrinsic quality factor was $Q_{\text{int}}=6.87\times 10^{4}$ and the external quality factor was $Q_{\text{ext}}=7.13\times 10^{4}$, so that the waveguide was approximately critically coupled. These results are as expected from our simulations. Note that there is a second dip in $|S_{\mathrm{ff}}|^{2}$ at lower frequency; this is likely an $\text{EH}_{11}$ mode, split by the rectangularity of the cavity.
In the initial configuration, the tunable mode had a slightly lower intrinsic quality factor of $5.6\times 10^{4}$, as expected, and a frequency of $2.8522\,\text{GHz}$. Each $1/4$ turn rotation of the tuning screws pushed the tuning plate inward by $0.20\,\mathrm{mm}$. As expected, this increased the tunable mode frequency by about $1\,\mathrm{MHz}$ at each step, as shown in Fig. 10a. We conservatively stopped after one full turn, as the tunable mode frequency became higher than that of the fixed mode. Throughout this procedure, the tunable mode’s quality factor, the coupling strengths of the waveguides, and all properties of the fixed mode remained essentially unchanged.
@@ 5.2 Suppression of Cross-Coupling
This measurement demonstrates that a cross-coupling $\chi\sim 10^{-4}$ can be achieved in our prototype cavity, corresponding to 80 dB of noise reduction. Even further reduction may be possible, but would not be visible in our measurement setup given the noise floor of the network analyzer.
@@ 6.1 Related Work and Future Directions
The modern wave of interest in heterodyne axion detection began in 2020 [9, 35, 10], and for the past several years, several experimental groups have been developing prototypes. The SHADE collaboration at Fermilab’s SQMS has characterized noise in a driven 9-cell superconducting cavity [12], and developed single cell cavities with a small, tunable splitting between the $\mathrm{TM}_{020}$ and $\mathrm{TE}_{011}$ mode [28]. The group also performed a sensitive search for dark photon dark matter using an unloaded superconducting cavity with fixed frequency [21]. The SHANHE collaboration at Peking University extended the latter result by demonstrating mechanical frequency tuning at cryogenic temperatures [49, 56], and is developing superconducting cavities optimized for heterodyne axion detection. The Australian UPLOAD collaboration performed an early search using the frequency technique [^3] The frequency technique was discussed in Sec. 2. The sensitivity projections for the frequency technique in Ref. [51] are weaker than projections for the power technique, but it was argued in Ref. [54] that the two methods have equivalent sensitivity if the dominant noise source is the noise temperature of the amplifier. The extent to which this holds will depend on the axion mass and the details of the experimental setup. [52], and another search using the power technique which demonstrated a $\sim 10\%$ tuning range [50]. These experiments were performed using non-superconducting cavities, but the group is developing superconducting cavities in collaboration with Jefferson Lab. Finally, the CERN Quantum Technology Initiative, in affiliation with the Physics Beyond Colliders group, has recently begun developing superconducting cavities for heterodyne axion detection with non-mechanical tuning mechanisms [2].
@@ 6.2 Outlook for a Superconducting Prototype
Our prototype has been designed from the ground up to be optimized for heterodyne axion detection. Its high signal form factor $C_{\text{sig}}\simeq 0.9$ compares favorably with the values $C_{\text{sig}}\simeq 0.5$ and $C_{\text{sig}}\simeq 0.2$ achievable in more traditional cylindrical or accelerator cavity designs, respectively, and could yield a substantially higher signal power due to the scaling $P_{\text{sig}}\propto C_{\text{sig}}^{2}$. Even more radical designs with toroidal cavity geometries have been proposed [35, 16], though it is less clear how to fabricate and tune them.
@@
Leakage noise measurements could be carried out in an existing cryogenic test facility, and would inform other aspects of the superconducting design. For instance, typical niobium cavities have a thickness of a few mm, while our current design has endplates of $20\,\mathrm{mm}$ thickness. Thicker walls of niobium would certainly be feasible and would greatly reduce vibrations, but may make thermal management more difficult. Moreover, stiffening the tuning membrane might lower the tuning range. The appropriate tradeoff can be determined by performing in situ measurements.
In an experimental run, one would need to calibrate the axion signal. As with other axion detection concepts, in the heterodyne approach it is not practical to insert a copy of the axion-induced effective current $\mathbf{J}_{\mathrm{eff}}$. However, the signal power is fully determined by the frequencies, quality factors, and external couplings of the cavity modes, which can be extracted from scattering measurements (as discussed in Sec. 5), and the signal form factor $C_{\mathrm{sig}}$ in Eq. (5). This form factor can be found numerically, as we have discussed, or directly measured by mapping out the modes using the “bead pull” method, standard in axion experiments.
@@
To estimate the potential sensitivity, we note that for a thermal noise limited heterodyne axion experiment with amplifier noise at the standard quantum limit, the signal to noise ratio was shown in Ref. [9] to be
SNR $\sim\frac{\rho_{\mathrm{DM}}\,V}{m_{a}\,\omega_{1}}\,(g_{a\gamma\gamma}\,C_{\text{sig}}\,B_{0})^{2}\sqrt{\frac{Q_{a}\,Q_{\text{int}}\,t_{e}}{T}}$ (12)
@@
$=5\times\left(\frac{g_{a\gamma\gamma}}{10^{-15}\,\text{G}\text{e\kern-0.6458ptV}^{-1}}\right)^{2}\frac{\rho_{\mathrm{DM}}}{0.4\,\text{G}\text{e\kern-0.6458ptV}/\mathrm{cm}^{3}}\frac{10^{-8}\,\text{e\kern-0.6458ptV}}{m_{a}}\left(\frac{Q_{a}}{10^{6}}\right)^{1/2}$
$\qquad\times\frac{V}{0.12\,\mathrm{m}^{3}}\frac{2\pi\,(2.85\,\mathrm{GHz})}{\omega_{1}}\left(\frac{C_{\text{sig}}}{0.9}\right)^{2}$
$\qquad\times\left(\frac{B_{0}}{0.03\,\mathrm{T}}\right)^{2}\left(\frac{Q_{\text{int}}}{3.9\times 10^{10}}\frac{t_{e}}{1\,\mathrm{week}}\frac{2\,\mathrm{K}}{T}\right)^{1/2}.$ (13)
@@
This estimate of the SNR holds up to an order-one factor which depends on details such as the precise shape of the axion frequency spectrum, and assumes the frequency scan is performed uniformly in $\log(m_{a})$. In the three lines of the second equality, we plug in standard parameters for dark matter at a representative mass and coupling, parameters which are already realized in our prototype cavity design, and parameters which would be reasonable in a future superconducting version of the cavity, as will be discussed in further detail below. Note that Eq. (12) implicitly assumes that the readout waveguide is optimally overcoupled by a factor of $T/\omega_{1}\simeq 15$, which optimizes the scan rate.
@@
For the cavity, we assume cooling to $T=2\,\mathrm{K}$, which is standard for a helium cryostat. In our prototype design, our simulations indicated that a tuning range of $10\,\mathrm{MHz}$ is both mechanically feasible and preserves the cavity mode properties. Only about half this tuning range was realized in the prototype cavity, since we could only push the tuning membrane inward. We assume that the motorized tuning mechanism for the superconducting cavity can both push and pull, realizing the full potential tuning range. It is simplest to estimate the sensitivity at high frequencies, where leakage noise is automatically suppressed, so we suppose the mode frequency difference is scanned from $1\,\mathrm{MHz}$ to $11\,\mathrm{MHz}$.
@@
As for the intrinsic quality factor $Q_{\text{int}}$, we have seen that a pure copper prototype cavity would have $Q_{\text{int}}=7\times 10^{4}$ and a surface resistance $R_{s}=\sqrt{\omega_{0}/2\sigma}=1.4\times 10^{-2}\,\Omega$, where $\sigma$ is the standard conductivity of room temperature copper. The surface resistance of niobium depends on frequency, temperature, and the manufacturing process. It can be decomposed as $R_{s}=R_{\text{BCS}}(f,T)+R_{\text{res}}(f)$, where the BCS contribution is proportional to $f^{2}$ and the residual resistance is due to impurities, defects, and trapped magnetic flux [32]. For the LCLS-II project, commercial vendors use nitrogen doping to reliably produce superconducting cavities with $R_{\text{BCS}}(1.3\,\mathrm{GHz},2\,\mathrm{K})=4.5\,\mathrm{n}\Omega$ and $R_{\text{res}}=2.5\,\mathrm{n}\Omega$ [29]. We assume the residual resistance is dominated by trapped flux, so that $R_{\text{res}}\propto f^{1/2}$ [32, 20]. This yields $R_{s}=25\,\mathrm{n}\Omega$ at the prototype cavity frequency, which gives a projected quality factor $Q_{\text{int}}=G/R_{s}=3.9\times 10^{10}$.
Though surface resistances in this range are routinely achieved on production-grade SRF cavities, it may seem challenging to achieve these values on complex surfaces, e.g. with grooves machined out of a thick niobium sheet. However, similar production techniques have already been demonstrated for HL-LHC crab cavities, with very good results [42]. We therefore believe there is no fundamental obstacle for obtaining state-of-the art results on machined surfaces, provided proper surface treatments like electropolishing and possibly nitrogen doping or infusion [25] are adopted, possibly with some development work.
In a superconducting implementation, the rectangular features in our cavity design could lead to multipacting, which could be alleviated by rounding the corners of the fins. Since addressing this question requires dedicated simulations, we defer a detailed study to future work.
@@
In Fig. 11, we show the projected reach of a superconducting version of our prototype at $\text{SNR}=2$. The “test run” assumes a total scan time of 1 week and an input power $P_{\mathrm{in}}=0.1\,\mathrm{W}$, which is within the reach of a standard low-noise oscillator. In this regime leakage noise would almost certainly be negligible, given the low values of $\chi$ and $\eta$ in our design, the small input power, and the $\gtrsim\mathrm{MHz}$ axion frequency. Note that, as discussed above, the signal mode should be highly overcoupled to the readout waveguide. In the optimal case this would reduce its quality factor to $Q_{1}\simeq 2.5\times 10^{9}$, which both decreases the time required for the signal to ring up, and simplifies the stabilization of the mode frequencies.
The “full run” curve assumes the same total scan time, but a loaded mode with typical magnetic field $B_{0}=0.03\,\mathrm{T}$. Maintaining this field requires an input power $P_{\mathrm{in}}=\omega_{0}B_{0}^{2}V/2Q_{\mathrm{int}}=20\,\mathrm{W}$, which is well within the capacity of a vertical test stand. At this input power, Table 2 implies a maximum magnetic field of $0.2\,\mathrm{T}$ and electric field of $60\,\mathrm{MV}/\mathrm{m}$ at the cavity walls, due to the field enhancement at the endplate fins. At these fields, proper surface treatment will be required to control field emission.
Finally, the “double scale” curve assumes a scan time of 1 year and a direct scale-up of the cavity while maintaining $B_{0}=0.03\,\mathrm{T}$. This automatically halves the mode frequencies, multiplies the volume by $8$, and increases the quality factor to $2.4\times 10^{11}$.